# Adversarially Robust Out-of-Distribution Detection Using Lyapunov-Stabilized Embeddings

### v1

@@ 1 Introduction
OOD detection is a safety-critical task that is crucial for deploying models in the real world. In this task, training is limited to in-distribution (ID) data, while the inference task involves identifying OOD samples, i.e., samples that deviate from the ID data (16; 17). Recent advancements have demonstrated impressive performance gains across various detection benchmarks (18; 19; 20; 21). However, a significant challenge arises concerning the robustness of OOD detectors against adversarial attacks. An adversarial attack on a detector involves introducing minor perturbations to test samples, causing the detector to predict OOD as ID samples or vice versa. Yet, a robust OOD detector is imperative, especially in scenarios like medical diagnostics and autonomous driving (22; 23; 24; 25; 26).
Recently, several approaches have sought to address this challenge by first demonstrating that relying solely on ID data is insufficient for building adversarially robust detectors (27; 28; 29; 23; 30; 26; 31; 32; 33; 34; 35; 36; 37; 38). Consequently, new methods propose incorporating copious amounts of auxiliary OOD data in conjunction with adversarial training to improve the detector’s robustness. While effective, a significant gap remains between detector performance on clean data and their robustness against adversarial attacks (see Figure 1, Tables 1, 2(a), and 2(b)).
@@ 3 Related Work
OOD Detection Methods. Existing OOD detection methods can be broadly categorized into post-hoc and training-based approaches. Post-hoc methods involve training a classifier on ID data and subsequently using statistics from the classifier’s outputs or intermediate representations to identify OOD samples. For instance, Hendrycks et al. (52) propose using the maximum softmax probability distributions (MSP) as a metric. The MD method (60) leverages the Mahalanobis Distance in the feature space, and OpenMax (61) recalibrates classification probabilities to improve OOD detection. Training-based methods, on the other hand, modify the training process to enhance OOD detection capabilities. Such modifications can include defining additional loss functions, employing data augmentation techniques, or incorporating auxiliary networks. Examples of training-based methods designed for standard OOD detection include VOS (39), DHM (19), CATEX (62), and CSI (63). On the other hand, ATOM (30), ALOE (28), ATD (34), and RODEO (37) have been developed specifically for robust OOD detection. For detailed descriptions of these methods, please refer to Appendix A1.
@@ 4 Proposed Method
Overview of AROS. AROS ensures that perturbed input samples remain close to their non-perturbed counterparts in the feature space by leveraging the Lyapunov stability theorem (71; 72; 73; 74). By using a NODE, we consider the model as a dynamical system and design it so that ID and OOD samples converge to distinct stable equilibrium points of that system. This approach prevents significant deviations in the output when adversarial perturbations are applied. However, since OOD data is unavailable, we craft fake OOD samples in the embedding space by estimating the boundaries of the ID distribution and sampling from the corresponding low-likelihood regimes. To further avoid any misprediction between OOD and ID data caused by perturbations, we maximize the distance between their equilibrium points by leveraging an orthogonal binary layer for classification. In the following sections, we will thoroughly explain each proposed component, highlighting the benefits of our approach.
@@ 4.1 Fake Embedding Crafting Strategy
We then select fake embeddings $r$ from the feature space corresponding to class $j$, where $j=1,\dots,K$, such that $r\sim\mathcal{N}(\hat{\mu}_{j},\hat{\Sigma}_{j})$ satisfies:
$\quad\quad\quad\quad\quad\quad p(r\mid y=j)=\frac{1}{(2\pi)^{d/2}|\hat{\Sigma}_{j}|^{1/2}}\exp\left(-\frac{1}{2}(r-\hat{\mu}_{j})^{T}\hat{\Sigma}_{j}^{-1}(r-\hat{\mu}_{j})\right)<\beta,\quad\quad\quad\quad(1)$
where we choose $\beta$ to be very small (e.g., 0.001) (Figure 2B); we also conduct an ablation study for different values of $\beta$ (see Appendix A3). Here, $d$ is the dimensionality of the feature vectors $r$, and $j=1,\dots,K$. The terms $\hat{\mu}_{j}$ and $\hat{\Sigma}_{j}$ represent the mean vector and covariance matrix of the $j$-th class of ID training samples in feature space, respectively:
$\quad\quad\quad\quad\quad\hat{\mu}_{j}=\frac{1}{n_{j}}\sum_{i:y_{i}=j}f_{\theta}(x_{i}),\quad\hat{\Sigma}_{j}=\frac{1}{n_{j}-1}\sum_{i:y_{i}=j}(f_{\theta}(x_{i})-\hat{\mu}_{j})(f_{\theta}(x_{i})-\hat{\mu}_{j})^{T},\quad\quad(2)$
where $n_{j}$ is the number of samples in class $j$. By sampling equally across each class of $\mathcal{D}^{\text{in}}_{\text{train}}$, we generate a set of synthetic, “fake” OOD embeddings (Figure 2C), denoted as $r_{\text{OOD}}$. We then construct a balanced training set by taking the union of the embeddings of ID samples and the OOD embeddings, defining it as: $X_{\text{train}}=\left\{f_{\theta}(\mathcal{D}^{\text{in}}_{\text{train}})\cup r_{\text{OOD}}\right\}.$ We define the labels $y$ for this set as 0 for ID and 1 for fake OOD embeddings. Lastly, at test-time we use this system to then predict the OOD score (Figure 2D).
@@ 4.3 Orthogonal Binary Layer and Training Step
(ResNet) (WideResNet) (CLIP) (ResNet) (DenseNet) (WideResNet) (WideResNet) (CLIP) (WideResNet)
@@ 5 Experiments
Here we present empirical evidence to validate the effectiveness of our method under various setups, including adversarial attacks, corrupted inputs (non-adversarial perturbations), and clean inputs (non-perturbed scenarios).
First, we adversarially train a classifier on ID data and then use it to map the data into a robust embedding space. A Gaussian distribution is fitted around these embeddings, and low-likelihood regions of the distribution are sampled to create fake OOD data as a proxy for OOD test samples. We then demonstrate that time invariance, which establishes that the NODE’s behavior does not explicitly depend on time, leads to more stable behavior of the detector under adversarial attacks (see Section 5). Consequently, we leverage Lyapunov stability regularization under a time-invariant assumption for training. However, a potential challenge arises when ID and OOD equilibrium points are located near each other. As a remedy, we introduce an orthogonal binary layer (86; 92; 93) that increases the separation between ID and OOD equilibrium points, thereby expanding the stability region around each point. This enhances the model’s robustness against shifting adversarial samples from OOD to ID and vice versa. Finally, we use the orthogonal binary layer’s confidence output as the OOD score during inference.
@@ 6 Ablation Study
AROS Components. To verify the effectiveness of AROS, we conducted ablation studies across various datasets. The corresponding results are presented in Table 4. In each experiment, individual components were replaced with alternative ones, while the remaining elements were held constant. In Config A, we ignored the designed loss function $\mathcal{L}_{\text{SL}}$ and instead utilized the cross-entropy loss function $\mathcal{L}_{\text{CE}}$ for binary classification. Config B represents the scenario in which we train the classifier in the first step without adversarial training on the ID data, instead using standard training. This reduces robustness as $f_{\theta}$ becomes more susceptible to perturbations within ID classes, ultimately making the final detector more vulnerable to attacks. In Config C, the orthogonal binary layer was replaced with a regular binary layer. In Config D, rather than estimating the ID distribution and sampling OOD data in the embedding space, we substituted this process by creating random noise in the embedding space as fake OOD data. This removes the conditioning of the fake OOD distribution on the ID data and, as a result, makes them unrelated. This is in line with previous works that have shown that related and nearby auxiliary OOD samples are more useful (43; 24). Config E represents our default pipeline. Finally, in Config F, we extended AROS by augmenting the fake OOD embedding data with additional OOD images (i.e., Food-101 (107)) alongside the proposed fake OOD strategy. Specifically, we transformed these additional OOD images into the embedding space using $f_{\theta}$ and combined them with the crafted fake embeddings, which led to enhanced performance.
@@ 7 Conclusions
Adversarially Robust OOD Detection Methods. Previous adversarially robust methods for OOD detection have built on the outlier exposure [27] technique and conducted adversarial training on a union of exposed outliers and ID samples. These methods primarily aim to enhance the outlier exposure technique by improving the diversity of the extra OOD image dataset. To tackle this, ALOE leverages an auxiliary OOD dataset and perturbs the samples to adversarially maximize the KL-divergence between the model’s output and a uniform distribution. It is important to note that the adversarial evaluation in their pipeline was less rigorous than that in our approach, due primarily to the limited budget allocated for adversarial attacks. ATOM, in contrast, selectively samples informative outliers rather than using random outliers. The adversarial evaluation of ATOM is also not entirely standard, as it exclusively targets OOD test samples. However, a more comprehensive approach would involve attacking both ID and OOD test samples to ensure enhanced robustness. Meanwhile, ATD employs a GAN to generate auxiliary OOD images for adversarial training, instead of relying on an external dataset. RODEO aims to demonstrate that enhancing the diversity of auxiliary OOD images, while maintaining their stylistic and semantic alignment with ID samples, will improve robustness. Consequently, they utilize a pretrained CLIP model [108] and a diffusion model for OOD image synthesis.
@@
Standard OOD Detection Methods. CSI enhances the outlier detection task by building on standard contrastive learning through the introduction of ’distributionally-shifted augmentations’—transformations that encourage the model to treat augmented versions of a sample as OOD. This approach enables the model to learn representations that more effectively distinguish between ID and OOD samples. The paper also proposes a detection score based on the contrastive features learned through this training scheme, which demonstrates effectiveness across various OOD detection scenarios, including one-class, multi-class, and labeled multi-class settings. DHM involves modeling the joint density of data and labels in a single forward pass. By factorizing this joint density into three sources of uncertainty (aleatoric, distributional, and parametric), DHMs aim to distinguish in-distribution samples from OOD samples. To achieve computational efficiency and scalability, the method employs weight normalization during training and utilizes normalizing flows (NFs) to model the probability distribution of the features. The key idea is to use bi-Lipschitz continuous mappings, enabled by spectral normalization, which allows the use of state-of-the-art deep neural networks for learning expressive and geometry-preserving representations of data. VOS uses virtual outliers to regularize the model’s decision boundary and improve its ability to distinguish between ID and OOD data. The framework includes an unknown-aware training objective that uses contrastive learning to shape the uncertainty surface between the known data and the synthesized outliers. This method is effective for both object detection and image classification. CATEX use of two hierarchical contexts—perceptual and spurious—to describe category boundaries more precisely through automatic prompt tuning in vision-language models like CLIP. The perceptual context distinguishes between different categories (e.g., cats vs. apples), while the spurious context helps identify samples that are similar but not in-distribution (e.g., distinguishing cats from panthers). This hierarchical structure helps create more precise category boundaries.
@@
(d) Table 7a: Comparison of AROS and AdvRAD under clean and $\text{PGD}^{1000}(l_{\infty})$ evaluation, measured by AUROC (%). The table cells denote results in the ‘ Clean/$\text{PGD}^{1000}$‘ format.
@@
A3.1.2 AROS Under AutoAttack (AA) and Adaptive AA
Additionally, we evaluate the robustness of AROS under different perturbation norms by substituting the $l_{\infty}$ norm with the $l_{2}$ norm in $\text{PGD}^{1000}$, setting $\epsilon=\frac{128}{255}$. The outcomes of this evaluation are presented in Table 4(e). Furthermore, the performance of AROS under AutoAttack (AA) and adaptive AA is included in Table 4(f).
(e) Table 8a: Evaluation of AROS under $\text{PGD}^{1000}(l_{2})$, where $\epsilon=128/255$ is used for low-resolution datasets and $\epsilon=64/255$ for high-resolution datasets. The results are measured by AUROC (%).
@@
(f) Table 9a: Evaluation of AROS under AutoAttack and Adaptive AA. The results are measured by AUROC (%).
@@
A3.1.3 Architecture comparison & Enhancing Clean Performance with Transfer Learning
Additionally, we investigate the influence of replacing our default backbone architecture, WideResNet, with alternative architectures (Table 4(g)). We also explore transfer learning techniques to enhance clean performance by leveraging robust pre-trained classifiers.
Lastly, while AROS demonstrates a significant improvement of up to 40% in adversarial robustness, it shows a performance gap of approximately 10% when compared to state-of-the-art clean detection methods. Although this trade-off between robustness and clean performance is well-documented in the literature [5, 105, 44], our aim is to enhance clean performance. A promising approach would involve enhancing clean performance while preserving robustness by leveraging transfer learning through distillation from a large, robust pretrained model, rather than training the classifier from scratch, as is currently done in the pipeline. Specifically, by utilizing adversarially pretrained classifiers on ImageNet [116], we aim to improve our clean performance. Results in lower part of Table 4(g) indicate that leveraging such pretrained models can improve clean performance by 6%. Moreover, we also consider using different architectures trained from scratch to further demonstrate the robustness of AROS across various backbones.
(g) Table 10a: Ablation study on different backbone architectures & Transfer learning. Results are reported under clean and $\text{PGD}^{1000}(l_{\infty})$ evaluations, measured by AUROC (%). Each table cell presents results in the ‘ Clean/$\text{PGD}^{1000}$‘ format.
@@
A3.2 Extended Ablation Study
Further ablation studies explore the sensitivity of AROS to various hyperparameters. The effect of the threshold parameter $\beta$, which is used to select fake OOD embeddings, is presented in Table 4(h), while the impact of varying $\gamma_{2}$ and $\gamma_{3}$ is detailed in Table 4(i).
(h) Table 11a: Ablation study on the $\beta$ hyperparameter. Results are reported under clean and $\text{PGD}^{1000}(l_{\infty})$ evaluations, measured by AUROC (%). Each table cell presents results in the ‘ Clean/$\text{PGD}^{1000}$‘ format.

### v2

@@ 1 Introduction
OOD detection is a safety-critical task that is crucial for deploying models in the real world. In this task, training is limited to in-distribution (ID) data, while the inference task involves identifying OOD samples, i.e., samples that deviate from the ID data (16; 17). Recent advancements have demonstrated impressive performance gains across various detection benchmarks (18; 19; 20; 21). However, a significant challenge arises concerning the robustness of OOD detectors against adversarial attacks. An adversarial attack on a detector involves introducing minor perturbations to test samples, causing the detector to predict OOD as ID samples or vice versa. Yet, a robust OOD detector is imperative, especially in scenarios like medical diagnostics and autonomous driving (22; 23; 24; 25; 26). Recently, several approaches have sought to address this challenge by first demonstrating that relying solely on ID data is insufficient for building adversarially robust detectors (27; 28; 29; 23; 30; 26; 31; 32; 33; 34; 35; 36; 37; 38). Consequently, new methods propose incorporating copious amounts of auxiliary OOD data in conjunction with adversarial training to improve the detector’s robustness. While effective, a significant gap remains between detector performance on clean data and their robustness against adversarial attacks (see Figure 1, Tables 1, 2(a), and 2(b)).
@@ 3 Related Work
OOD Detection Methods. Existing OOD detection methods can be broadly categorized into post-hoc and training-based approaches. Post-hoc methods involve training a classifier on ID data and subsequently using statistics from the classifier’s outputs or intermediate representations to identify OOD samples. For instance, Hendrycks et al. (52) propose using the maximum softmax probability distributions (MSP) as a metric. The MD method (60) leverages the Mahalanobis Distance in the feature space, and OpenMax (61) recalibrates classification probabilities to improve OOD detection. Training-based methods, modify the training process to enhance OOD detection capabilities. Such modifications can include defining additional loss functions, employing data augmentation techniques, or incorporating auxiliary networks. Examples of training-based methods designed for standard OOD detection include VOS (39), DHM (19), CATEX (62), and CSI (63). On the other hand, ATOM (30), ALOE (28), ATD (34), and RODEO (37) have been developed specifically for robust detection. For detailed descriptions of these methods, please refer to Appendix A1.
@@ 4 Proposed Method
Overview of AROS. AROS ensures that perturbed input samples remain close to their non-perturbed counterparts in the feature space by leveraging the Lyapunov stability theorem (71; 72; 73; 74). By using a NODE, we consider the model as a dynamical system and design it so that ID and OOD samples converge to distinct stable equilibrium points of that system. This approach prevents significant deviations in the output when adversarial perturbations are applied. However, since OOD data is unavailable, we craft fake OOD samples in the embedding space by estimating the boundaries of the ID distribution and sampling from the corresponding low-likelihood regimes. To further avoid any misprediction between OOD and ID data caused by perturbations, we maximize the distance between their equilibrium points by leveraging an orthogonal binary layer for classification. In the following, we will thoroughly explain each proposed component, highlighting the benefits of AROS.
@@ 4.1 Fake Embedding Crafting Strategy
We then select fake embeddings $r$ from the feature space corresponding to class $j$ such that $r\sim\mathcal{N}(\hat{\mu}_{j},\hat{\Sigma}_{j})$ satisfies:
$\quad\quad\quad\quad\quad\quad\quad\quad\frac{1}{(2\pi)^{d/2}|\hat{\Sigma}_{j}|^{1/2}}\exp\left(-\frac{1}{2}(r-\hat{\mu}_{j})^{T}\hat{\Sigma}_{j}^{-1}(r-\hat{\mu}_{j})\right)<\beta,\quad\quad\quad\quad\quad\quad\quad(1)\quad$ where, $\beta$ serves as a likelihood threshold, and we set that to a small value (e.g., 0.001) (Figure 2B). Additionally, we conduct an ablation study to evaluate the impact of different values of $\beta$ and discuss practical considerations (see Appendix A3.3).Our comprehensive ablation experiments demonstrate the consistent performance of AROS across varying $\beta$ values.
Here, $d$ is the dimensionality of the feature vectors $r$, and $j=1,\dots,K$. The terms $\hat{\mu}_{j}$ and $\hat{\Sigma}_{j}$ represent the mean vector and covariance matrix of the $j$-th class of ID training samples in feature space, respectively:
$\quad\quad\quad\quad\quad\hat{\mu}_{j}=\frac{1}{n_{j}}\sum_{i:y_{i}=j}f_{\theta}(x_{i}),\quad\hat{\Sigma}_{j}=\frac{1}{n_{j}-1}\sum_{i:y_{i}=j}(f_{\theta}(x_{i})-\hat{\mu}_{j})(f_{\theta}(x_{i})-\hat{\mu}_{j})^{T},\quad\quad(2)$ where $n_{j}$ is the number of samples in class $j$. By sampling equally across each class of $\mathcal{D}^{\text{in}}_{\text{train}}$, we generate a set of synthetic, “fake” OOD embeddings (Figure 2C), denoted as $r_{\text{OOD}}$. We then construct a balanced training set by taking the union of the embeddings of ID samples and the OOD embeddings, defining it as: $X_{\text{train}}=\left\{f_{\theta}(\mathcal{D}^{\text{in}}_{\text{train}})\cup r_{\text{OOD}}\right\}.$ We define the labels $y$ for this set as 0 for ID and 1 for fake OOD embeddings.
@@ 4.3 Orthogonal Binary Layer and Training Step
(ResNet) (WideResNet) (CLIP-ViT) (ResNet) (DenseNet) (WideResNet) (WideResNet) (CLIP-ViT) (WideResNet)
@@ 5 Experiments
Here we present empirical evidence to validate the effectiveness of our method under various setups, including adversarial attacks, corrupted inputs (non-adversarial perturbations), and clean inputs (non-perturbed scenarios). We note that the backbone architecture for the methods considered is the same as described in Table 1
First, we adversarially train a classifier on ID data and then use it to map the data into a robust embedding space. A Gaussian distribution is fitted around these embeddings, and low-likelihood regions of the distribution are sampled to create fake OOD data as a proxy for OOD test samples. We then demonstrate that time invariance, which establishes that the NODE’s behavior does not explicitly depend on time, leads to more stable behavior of the detector under adversarial attacks (see Section 4.3). Consequently, we leverage Lyapunov stability regularization under a time-invariant assumption for training. However, a potential challenge arises when ID and OOD equilibrium points are located near each other. As a remedy, we introduce an orthogonal binary layer (86; 92; 93) that enhances the separation between ID and OOD data by increasing the distance between their neighborhoods of Lyapunov-stable equilibrium. This enhances the model’s robustness against shifting adversarial samples from OOD to ID and vice versa. Finally, we use the orthogonal binary layer’s confidence output as the OOD score during inference.
@@ 6 Ablation Study
AROS Components. To verify the effectiveness of AROS, we conducted ablation studies across various datasets. The corresponding results are presented in Table 4. In each experiment, individual components were replaced with alternative ones, while the remaining elements were held constant. In Config A, we ignored the designed loss function $\mathcal{L}_{\text{SL}}$ and instead utilized the cross-entropy loss function $\mathcal{L}_{\text{CE}}$ for binary classification. Config B represents the scenario in which we train the classifier in the first step without adversarial training on the ID data, instead using standard training. This reduces robustness as $f_{\theta}$ becomes more susceptible to perturbations within ID classes, ultimately making the final detector more vulnerable to attacks. In Config C, the orthogonal binary layer was replaced with a regular binary layer. In Config D, rather than estimating the ID distribution and sampling OOD data in the embedding space, we substituted this process by creating random Gaussian noise in the embedding space as fake OOD data. This removes the conditioning of the fake OOD distribution on the ID data and, as a result, makes them unrelated. This is in line with previous works that have shown that related and nearby auxiliary OOD samples are more useful (43; 24). Config E represents our default pipeline. Finally, in Config F, we extended AROS by augmenting the fake OOD embedding data with additional OOD images (i.e., Food-101 (107)) alongside the proposed fake OOD strategy. Specifically, we transformed these additional OOD images into the embedding space using $f_{\theta}$ and combined them with the crafted fake embeddings, which led to enhanced performance.
@@ 7 Conclusions
Adversarially Robust OOD Detection Methods. Previous adversarially robust methods for OOD detection have built on the outlier exposure [27] technique and conducted adversarial training on a union of exposed outliers and ID samples. These methods primarily aim to enhance the outlier exposure technique by improving the diversity of the extra OOD image dataset. To tackle this, ALOE leverages an auxiliary OOD dataset and perturbs the samples to adversarially maximize the KL-divergence between the model’s output and a uniform distribution. It is important to note that the adversarial evaluation in their pipeline was less rigorous than that in our approach, due primarily to the limited budget allocated for adversarial attacks. ATOM, in contrast, selectively samples informative outliers rather than using random outliers. The adversarial evaluation of ATOM is also not entirely standard, as it exclusively targets OOD test samples. However, a more comprehensive approach would involve attacking both ID and OOD test samples to ensure enhanced robustness. Meanwhile, ATD employs a GAN to generate auxiliary OOD images for adversarial training, instead of relying on an external dataset. RODEO aims to demonstrate that enhancing the diversity of auxiliary OOD images, while maintaining their stylistic and semantic alignment with ID samples, will improve robustness. Consequently, they utilize a pretrained CLIP model [108] and a diffusion model for OOD image synthesi[43, 109, 54, 110, 111, 112].
@@
Standard OOD Detection Methods. CSI enhances the outlier detection task by building on standard contrastive learning through the introduction of ’distributionally-shifted augmentations’—transformations that encourage the model to treat augmented versions of a sample as OOD. This approach enables the model to learn representations that more effectively distinguish between ID and OOD samples. The paper also proposes a detection score based on the contrastive features learned through this training scheme, which demonstrates effectiveness across various OOD detection scenarios, including one-class, multi-class, and labeled multi-class settings. DHM involves modeling the joint density of data and labels in a single forward pass. By factorizing this joint density into three sources of uncertainty (aleatoric, distributional, and parametric), DHMs aim to distinguish in-distribution samples from OOD samples. To achieve computational efficiency and scalability, the method employs weight normalization during training and utilizes normalizing flows (NFs) to model the probability distribution of the features. The key idea is to use bi-Lipschitz continuous mappings, enabled by spectral normalization, which allows the use of state-of-the-art deep neural networks for learning expressive and geometry-preserving representations of data. VOS uses virtual outliers to regularize the model’s decision boundary and improve its ability to distinguish between ID and OOD data. The framework includes an unknown-aware training objective that uses contrastive learning to shape the uncertainty surface between the known data and the synthesized outliers. This method is effective for both object detection and image classification. CATEX use of two hierarchical contexts—perceptual and spurious—to describe category boundaries more precisely through automatic prompt tuning in vision-language models like CLIP. The perceptual context distinguishes between different categories (e.g., cats vs. apples), while the spurious context helps identify samples that are similar but not ID (e.g., distinguishing cats from panthers). This hierarchical structure helps create more precise category boundaries.
@@
(d) Table 7a: Comparison of AROS and AdvRAD under clean and $\text{PGD}^{1000}(l_{\infty})$ evaluation, measured by AUROC (%). The table cells denote results in the ‘ Clean/$\text{PGD}^{1000}$‘ format. The perturbation budget $\epsilon$ is set to $\frac{8}{255}$ for low-resolution datasets and $\frac{4}{255}$ for high-resolution datasets.
@@
A3.1.2 AROS Under Additional Adversarial Attacks
Figure 3: Visualization of clean and perturbed images from the CIFAR-10 dataset to illustrate the impact of perturbations on semantic content. The first row depicts clean images, while the second and third rows show images perturbed with $L_{\infty}$ norm of $\frac{8}{255}$ and $L_{2}$ norm of $\frac{128}{255}$, respectively. Despite the added perturbations, the semantic content of the images remains unchanged, demonstrating that robustness expectations from models under these perturbations are fair.
In this section, we further emphasize the robustness of AROS by evaluating its performance under different attacks. First, we conducted additional experiments considering different perturbation norms, specifically substituting the $l_{\infty}$ norm with the $l_{2}$ norm in $\text{PGD}^{1000}$, setting $\epsilon=\frac{128}{255}$. The outcomes of this evaluation are presented in Table 4(e). See Figure 3 for an illustration of some clean and perturbed samples.
Next, we evaluated the robustness of AROS under $\text{PGD}^{1000}(l_{\infty})$ with a higher perturbation budget (i.e., $\epsilon=\frac{16}{255}$), as shown in Table 4(f). Furthermore, the performance of AROS under AutoAttack (AA) and adaptive AA is reported in Table 4(g).
Additionally, evaluating a model against a variety of transfer-based adversarial attacks is crucial for understanding its robustness in real-world scenarios, where adversarial examples crafted for one model can successfully deceive others. These attacks simulate diverse and challenging conditions by targeting features such as generalization, input transformations, and model invariance, providing a comprehensive assessment of the model’s resilience. Such evaluations reveal potential vulnerabilities, measure performance under adversarial conditions, and offer insights into the model’s ability to generalize across different attack strategies.
Motivated by this, we evaluate our method, as well as DHM and RODEO, against several adversarial attacks, including DTA [120], DeCoWA [121], and SASD-WS [122]. We selected DHM and RODEO for comparison due to their strong performance in clean detection and robust detection, respectively. For these experiments, we utilized the implementation provided in the https://github.com/Trustworthy-AI-Group/TransferAttack repository. The results of this evaluation are presented in Table 4(h).
The consistent and superior performance of AROS against the aforementioned attacks underscores its effectiveness in adversarial scenarios.
(e) Table 8a: Evaluation of AROS under $\text{PGD}^{1000}(l_{2})$, where $\epsilon=\frac{128}{255}$ is used for low-resolution datasets and $\epsilon=\frac{64}{255}$ for high-resolution datasets. The results are measured by AUROC (%).
@@
(f) Table 8b: : Evaluation of AROS under $\text{PGD}^{1000}(l_{\infty})$, where $\epsilon=\frac{16}{255}$ is used for all datasets, including both high- and low-resolution ones. The results are reported in terms of AUROC (%).
Attack CIFAR-10 CIFAR-100 ImageNet-1k
CIFAR-100 SVHN LSUN CIFAR-10 SVHN LSUN Texture iNaturalist LSUN
$\text{PGD}^{1000}(l_{\infty})$ 70.3 78.4 72.6 58.2 63.9 60.4 60.5 66.8 62.7
(g) Table 9a: Evaluation of AROS under AutoAttack and Adaptive AA. The perturbation budget $\epsilon$ is set to $\frac{8}{255}$ for low-resolution datasets and $\frac{4}{255}$ for high-resolution datasets.
@@
(h) Table 9b: Comparison of AROS, DHM, and RODEO under Transfer-Based Attacks. The perturbation budget $\epsilon$ is set to $\frac{16}{255}$. Results compare AROS, DHM, and RODEO across DTA, DeCoWA, and SASD-WS attacks.
Attack Methods CIFAR-10 CIFAR-100 ImageNet-1k
CIFAR-100 SVHN LSUN CIFAR-10 SVHN LSUN Texture iNaturalist LSUN
AROS 84.7 86.4 88.2 71.5 76.6 71.7 76.2 83.5 75.9
DTA RODEO 61.6 68.5 89.3 51.1 63.1 70.9 61.1 60.6 59.2
DHM 72.9 78.3 78.9 75.4 79.1 74.1 55.9 59.2 57.6
AROS 83.8 85.2 86.8 69.6 77.2 69.4 72.6 79.5 74.8
DeCoWA RODEO 57.8 64.6 80.4 43.0 57.7 67.5 54.7 54.8 52.5
DHM 60.1 64.3 64.9 68.9 62.1 60.8 54.9 58.4 56.0
AROS 82.3 84.9 81.7 67.6 75.0 67.7 72.4 77.5 71.2
SASD-WS RODEO 49.5 53.4 79.4 44.7 56.5 66.1 53.2 46.9 47.0
DHM 64.6 69.3 61.3 56.3 67.8 59.7 53.1 51.3 49.2
A3.2 Detecting Backdoored Samples
We conducted an experiment to evaluate our method’s ability to detect clean samples (without triggers) as ID and poisoned samples as OOD. The results are presented in this section, with STRIP [123] included as a baseline method for identifying Trojaned samples.
It is important to highlight a key distinction in the context of OOD detection, which underpins our study. In traditional OOD detection scenarios, ID and OOD datasets typically differ at the semantic level (e.g., CIFAR-10 versus SVHN). However, detecting backdoored samples presents a unique challenge: the presence of triggers modifies clean images at the pixel level rather than the semantic level. Consequently, this task often hinges on texture-level differences rather than semantic distinctions, necessitating models designed specifically to leverage this inductive bias effectively.
Despite these challenges, our results demonstrate that AROS achieves strong performance in detecting poisoned samples. This underscores the versatility and effectiveness of AROS, even when applied to the nuanced problem of backdoor detection. The results are presented in Table 4(i).
(i) Table 10a: Effectiveness of the Proposed Method for Detecting Backdoor Attack Samples across CIFAR-10, CIFAR-100, and GTSRB datasets. The results are presented for different backdoor attack methods, demonstrating the performance of AROS and STRIP.
Method Backdoor Attack CIFAR-10 CIFAR-100 GTSRB
Badnets 80.3 67.5 72.8
AROS Wanet 62.7 58.9 54.4
SSBA 57.2 72.6 66.0
Badnets 79.2 86.0 87.1
STRIP Wanet 39.5 48.5 35.6
SSBA 36.4 68.5 64.1
A3.2.1 Architecture comparison & Enhancing Clean Performance with Transfer Learning
Additionally, we investigate the influence of replacing our default backbone architecture, WideResNet, with alternative architectures (Table 4(j)). We also explore transfer learning techniques to enhance clean performance by leveraging robust pre-trained classifiers.
Lastly, while AROS demonstrates a significant improvement of up to 40% in adversarial robustness, it shows a performance gap of approximately 10% when compared to state-of-the-art clean detection methods. Although this trade-off between robustness and clean performance is well-documented in the literature [5, 105, 44], our aim is to enhance clean performance. A promising approach would involve enhancing clean performance while preserving robustness by leveraging transfer learning through distillation from a large, robust pretrained model, rather than training the classifier from scratch, as is currently done in the pipeline. Specifically, by utilizing adversarially pretrained classifiers on ImageNet [124], we aim to improve our clean performance. Results in lower part of Table 4(j) indicate that leveraging such pretrained models can improve clean performance by 6%. Moreover, we also consider using different architectures trained from scratch to further demonstrate the robustness of AROS across various backbones.
(j) Table 11a: Ablation study on different backbone architectures & Transfer learning. Results are reported under clean and $\text{PGD}^{1000}(l_{\infty})$ evaluations, measured by AUROC (%). Each table cell presents results in the ‘ Clean/$\text{PGD}^{1000}$‘ format.
@@
A3.3 Hyperparameter Ablation Study
A3.3.1 Ablation Study on Hyperparameter of Fake Generation ($\beta$)
To craft fake samples, we fit a GMM to the embedding space of ID samples. The objective is to sample from the GMM such that the likelihood of the samples is low, ensuring they do not belong to the ID distribution and are located near its boundaries. This approach generates near-OOD samples, which are valuable for understanding the distribution manifold and improving the detector’s performance.
In practice, an encoder is used as a feature extractor to obtain embeddings of ID samples. A GMM is then fitted to these embeddings, and the likelihood of each training sample under the GMM is computed. The embeddings are subsequently sorted based on their likelihoods, and the $\beta$-th minimum likelihood of the training samples is used as a threshold. Random samples are drawn from the GMM, and their likelihoods are compared against this threshold. If a sample’s likelihood is lower than the threshold, it is retained as fake OOD; otherwise, it is discarded.
For instance, setting $\beta=0.1$ ensures that the crafted fake OOD samples have a likelihood of belonging to the ID distribution that is lower than 90% of the ID samples. Conversely, setting $\beta=0.5$ corresponds to randomly sampling from the ID distribution, rather than targeting low-likelihood regions, which leads to poor performance. An ablation study in our manuscript explores the effects of varying $\beta$ values on model performance.
The sensitivity of AROS to this hyperparameter is analyzed in detail. As shown in Table 4(k), AROS demonstrates consistent performance for small values of $\beta$. Our experimental results indicate that selecting $\beta$ values in the range $[0.0,0.1]$ achieves optimal performance, highlighting the robustness of AROS to changes in $\beta$.
(k) Table 12a: Ablation study on the $\beta$ hyperparameter. Results are reported under clean and $\text{PGD}^{1000}(l_{\infty})$ evaluations, measured by AUROC (%). Each table cell presents results in the ‘ Clean/$\text{PGD}^{1000}$‘ format.